OpenStudy (anonymous):

A compound is found to be 20 percent carbon, 1 percent hydrogen, and 12 percent nitrogen, with the rest being oxygen. What is the empirical formula of this compound? a. CHO b. C4H6O10 C. C2HNO5 d. Cannot determine from the information provided

OpenStudy (anonymous):

We are given percentages of each element comprising the compound so look at each percent as 1 gram. Then divide the number by the molar mass of each element. 20g Carbon/12.01g/mol Carbon = 1.66 mol 1g Hydrogen/1.01g/mol Hydrogen= 1.00 mol 12g Nitrogen/ 14.01g/mol Nitrogen= 0.86 mol 67g Oxygen/ 16.00g/mol Oxygen= 4.19 mol Dividing the number of moles of each element by the smallest number of moles (in this case Nitrogen, with 0.86 moles) would give us the empirical formula. Carbon= 1.66/0.86 =2 Hydrogen= 1.00/0.86 =1 Nitrogen= 0.86/0.86= 1 Oxygen= 4.19/0.86= 5 The number we have solved for in respect to each element is the amount of that element in the empirical formula. Therefore, our empirical formula is answer C: \[C _{2}HNO _{_{5}}\] Hope I helped!